According to the Taylor rule, if inflation is 8 percent and the GDP gap is 3 percent, what is the recommendation for the federal funds rate target?
1
Expert's answer
2020-10-08T12:07:37-0400
"Solution"

John Taylor proposed that monetary authorities use a rule for making discretionary decisions.

Taylor's rule states that; the federal funds rate target should be one-and-a- half times the inflation rate plus one-half times the GDP gap plus one.

That is,

"Federal\ fund\ rate=1.5(inflation\ rate)+0.5(GDP\ gap
)+1\\
 =1.5(8\%)+0.5(3\%)+1\\
=14.5\% (Which\ is\ the\ recommended\ federal\ funds\ rate\ target.)"